The probability that an industry is located in city A is 0.7; that it is located in city B is 0.4 and that it is located in A or B, or in both it is 0.8.

What is the probability that the industry is located:

a) In both cities?

b) In any of them?
Answer #1

a)

P(A or B) = P(A) + P(B ) - P(A and B)

0.8 = 0.7 + 0.4 - P(A and B)

Therefore,

P(both A and B) = 0.3

b)

P(Any of them) = P(A or B) = 0.8